OpenStudy (anonymous):

Can someone check to see if I'm right? Do the data in the table represent a direct variation or an inverse variation? Write an equation to model the data in the table. xl 1 l 3 l 3 l 7 l yl 5 l 15 l 20 l 35 l direct variation; y= 1/5x inverse variation; xy= 5 direct variation; y= 5x inverse variation; xy= 1/5 I think its direct variation; y=5x

OpenStudy (kendricklamar2014):

\[5*1 = 5\] \[5*3 = 15\] \[5*4 = 20\] \[5*7 = 35\] This proves that choice C is the answer :)